During the Covid-19 pandemic, the Ayurvedic medicines have an inelastic demand and electronic devices have an elastic demand. Imagine that technological advancement doubles the supply of both products (i.e. quantity supplied at each price is twice as it was earlier)


i. What will be the equilibrium price and quantity in each market?


ii. Which product experiences a larger change in price and which product experiences larger change in quantity. (9)


b. Do you agree with this statement, MC curve crosses the ATC curve at its minimum?


Why? (4)


c. How does an increase in interest rate affects household savings? Explain with the help of an example with graph.


1
Expert's answer
2021-03-26T11:53:49-0400

Solution:

i.). The linear function: y = Mx + b

Market 1: Ayurvedic medicines

Demand function (Qd) = 2P - 2

Supply function (Qs) = 2P + 2

At equilibrium: Qs = Qd

2P-2 = 2P+2

2P-2P = 2+2

P = 4

Q = 2(4) – 2= 8 – 2 = 6

Equilibrium Price = 4

Equilibrium Quantity = 6

Market 2: Electronic devices

Demand function (Qd) = 4P - 4

Supply function (Qs) = 4P + 4

At equilibrium: Qs = Qd

4P-4 = 4P+4

4P-4P = 4+4

P = 8

Q = 4(8) – 2= 30 – 2 = 28

Equilibrium Price = 8

Equilibrium Quantity = 28


ii). Electronic devices experience both a larger change in price and a larger change in quantity since it has an elastic demand.

b.). Yes, I agree with the statement.

The marginal cost always intersects the average total cost curve at its lowest point because the marginal cost of making the next unit of output will always affect the average total cost.


c.). Interest rates normally determine the amount of interest payments that savers will receive on their deposits. An increase in interest rates will make saving more attractive and should encourage saving. A drop in interest rates will reduce the benefits of saving and will tend to discourage saving.


This has been depicted by the below graph: